OpenStudy (anonymous):

The straight line with equation3/4x makes an acute angle theta with the x-axis. Find the value of sin2theta and cos2theta. how do i do this?

OpenStudy (anonymous):

Have you learned basic trig functions?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\tan \theta = \frac{ y }{ x } = \frac{ 3/4 }{ 1 }\]

OpenStudy (anonymous):

\[\theta = \tan^{-1} \left( \frac{ 3 }{ 4 } \right)\]

OpenStudy (anonymous):

Does this help, or do you need to solve it not using trig functions?

OpenStudy (anonymous):

thanks (:. this helped a lot. i think i understand now